# NewtonBench: Benchmarking Generalizable Scientific Law Discovery in LLM Agents

### v1

@@ Abstract
Large language models (LLMs) are emerging as powerful tools for scientific law discovery, a foundational challenge in AI-driven science. However, existing benchmarks for this task suffer from a fundamental methodological trilemma, forcing a trade-off between scientific relevance, scalability, and resistance to memorization. Furthermore, they oversimplify discovery as static function fitting, failing to capture the authentic scientific process of uncovering embedded laws through the interactive exploration of complex model systems. To address these critical gaps, we introduce NewtonBench, a benchmark comprising 324 scientific law discovery tasks across 12 physics domains. Our design mitigates the evaluation trilemma by using metaphysical shifts—systematic alterations of canonical laws—to generate a vast suite of problems that are scalable, scientifically relevant, and memorization-resistant. Moreover, we elevate the evaluation from static function fitting to interactive model discovery, requiring agents to experimentally probe simulated complex systems to uncover hidden principles. Our extensive evaluation of 11 state-of-the-art LLMs reveals a clear but fragile capability for discovery in frontier models: this ability degrades precipitously with increasing system complexity and exhibits extreme sensitivity to observational noise. Notably, we uncover a paradoxical effect of tool assistance: providing a code interpreter can hinder more capable models by inducing a premature shift from exploration to exploitation, causing them to satisfice on suboptimal solutions. These results demonstrate that robust, generalizable discovery in complex, interactive environments remains the core challenge for the future of automated science. By providing a scalable, robust, and scientifically authentic testbed, NewtonBench offers a crucial tool for measuring true progress and guiding the development of next-generation AI agents capable of genuine scientific discovery.
@@ 1 Introduction
Scientific law discovery, the process of distilling complex natural phenomena into concise, predictive mathematical principles, represents a cornerstone of human intellectual achievement, having consistently catalyzed paradigm shifts across the sciences (Langley et al., 1987; Popper, 2002). This endeavor demands a sophisticated interplay of abilities central to the scientific method: formulating hypotheses, methodical experimentation, and synthesizing empirical findings into universal laws (Peirce, 1958; Schmidt & Lipson, 2009).
@@
AI Feynman (Udrescu & Tegmark, 2020) 120 Function Partially Tunable Passive Observation
EmpiricalBench (Cranmer, 2023) 9 Function Fixed Passive Observation
@@
LLM-SR (Shojaee et al., 2025a) 4 Function Fixed Passive Observation
EvoSLD (Lin et al., 2025) 5 Function Fixed Passive Observation
@@
LSR-Transform (Shojaee et al., 2025b) 111 Function Fixed Passive Observation
@@
LSR-Synth (Shojaee et al., 2025b) 128 Function Fixed Passive Observation
PhysSymbol (Liu et al., 2025) 5,000 Function Tunable Passive Observation
@@
The rapid advancement of Large Language Models (LLMs) has brought this grand challenge into sharp focus (OpenAI, 2024; Google, 2025). Exhibiting remarkable emergent capabilities in mathematical reasoning (Ahn et al., 2024), logical reasoning (Yu et al., 2024), and agentic planning (Huang et al., 2024b), these models appear to possess foundational skills for scientific inquiry. This has sparked widespread curiosity (Armitage, 2025; Fang et al., 2025), crystallized by provocative questions such as whether an LLM could “rediscover Newton’s laws from its own intelligence(He, 2023). However, answering such questions requires moving beyond anecdotal observations. To foster and guide progress toward automated science, it is critical to develop a robust testbed to rigorously benchmark the true abilities and inherent limitations of current LLMs in this domain.
@@
While the trilemma highlights the trade-offs between existing paradigms, a more fundamental limitation underlies them all: they frame scientific discovery as a problem of static function discovery, where the goal is simply to fit a mathematical formula to variables in a tabular dataset. This static framing is a stark departure from the authentic scientific process, which relies on the interactive exploration of complex model systems (Cartwright, 1983; de Silva et al., 2020)—for instance, J.J. Thomson’s inference of the e/m ratio from cathode ray manipulation. Therefore, to meaningfully benchmark LLMs, evaluations must move beyond static function fitting and embrace the more scientifically authentic challenge of model discovery in interactive environments (Fang et al., 2025).
To address these limitations and facilitate a more rigorous evaluation, we introduce NewtonBench. Our design is built on two core principles, each targeting a fundamental shortcoming in prior work. First, to resolve the methodological trilemma, we introduce the metaphysical shift (or counterfactual shift)—a technique that systematically alters the mathematical structure of canonical physical laws, for instance, by modifying operators or exponents (see examples in Figure 1). This generates a vast suite of laws that are conceptually grounded but physically novel. This approach simultaneously achieves the scalability of synthetic laws, maintains the scientific relevance of real-world principles, and ensures resistance to memorization by creating problems that cannot be solved by recall, thus forcing models to reason from first principles.
@@
Furthermore, NewtonBench is designed to precisely probe the limits of an LLM’s scientific capabilities. It features two independent dimensions of difficulty control: the intrinsic complexity of the target law (tuned via the metaphysical shift) and the extrinsic complexity of the surrounding experimental system. This allows for a fine-grained analysis of a model’s breaking points. To isolate the challenge of discovery itself, we introduce an advanced setting where the LLM is provided with a code execution interface for tasks like numerical regression or hypothesis testing. By removing computational limitations as a primary bottleneck, this setup is designed to push models from being merely compute-bound to being truly discovery-bound, revealing the genuine frontiers of their scientific reasoning abilities.
@@ 2 Task Formalization
Figure 1: An illustration of core designs in NewtonBench: experimentation with model system, metaphysical shifts in physical laws from various domains, and agentic exploration settings.
@@ 3.1 Physical Law Curation
Our benchmark’s physical law curation process involves two primary stages: first, the collection of canonical seed laws, and second, the application of Metaphysical Shifts to generate novel variations. For the first stage, the selection of our seed laws is guided by three criteria: they must be foundational to their respective fields, span a diverse range of physics domains, and, critically, each must allow for dimensionally consistent mutations (often via adjusting physical constants).
The second stage involves applying the principle of Metaphysical Shifts (Wang & Song, 2025), which serves primarily to prevent agents from solving tasks by memorizing known laws. This principle is grounded in philosophical concepts regarding the contingency of natural laws and the existence of possible worlds [^1] This approach is analogous to the use of counterfactual reasoning in physics (Elwenspoek, 2012). (Lewis, 1986; Lange, 2008). In our context, a metaphysical shift is a systematic alteration of a canonical physical law, which effectively creates a novel yet physically plausible universe for the agent to discover in each task.
Operationally, these shifts are implemented as mutation operations (Koza, 1994) on the equation’s expression tree. These mutations can alter either the mathematical operators (e.g., changing an addition to a multiplication) or the values of numeric constants (e.g., modifying an exponent to change a quadratic relationship to a cubic one). A critical consequence of such alterations is the potential disruption of dimensional consistency. To preserve the physical coherence of the modified equation, this disruption is offset by systematically adjusting the dimensional units of the embedded physical constant. This necessity dictates that every target law in our benchmark includes at least one such constant to serve this compensatory role.
Three levels of equation difficulty are categorized based on the cumulative extent of mutations applied to a canonical law [^2] Consequently, equation difficulty arises from two factors: equation complexity and the level of distributional shift.. Easy laws are generated via 1–2 mutations from the original law. Medium laws are subsequently derived from Easy laws through 1–2 additional mutations, and Hard laws are similarly derived from Medium laws. We curated 108 shifted laws from 12 canonical laws under strict guidelines to ensure scientific plausibility, conducted by three domain‑expert co‑authors. The size of our shifted laws is comparable to existing synthetic benchmarks (Shojaee et al., 2025b). Appendix A.3.3 provides the complete set of laws, their mutated forms, and the curation guidelines.
@@ 3.3 Evaluation Metrics
To measure the data fidelity of a submitted law, we adopt the RMSLE metric. Unlike the metric of Normalized Mean Squared Error (NMSE) used in some prior studies (Shojaee et al., 2025b; Lin et al., 2025), RMSLE offers superior numerical stability and is better suited for physical quantities that may span multiple orders of magnitude due to its logarithmic scale. Given a dataset of $N$ points with ground-truth values $y_{i}$ and corresponding predictions $\hat{y}_{i}$ from the submitted law, RMSLE is calculated as (implementation details in Appendix A.4.2):
@@
GPT-4.1-mini 20.8   (±9.488) 2.8   (±2.268) 0.7   (±1.389) 10.4   (±2.660) 1.4   (±2.778) 0.0   (±0.000) 5.6   (±2.268) 0.0   (±0.000) 0.0   (±0.000) 4.6   (±1.260) 4.3363   (±0.358)   1.80
GPT-4.1 20.1   (±4.744) 1.4   (±1.604) 0.0   (±0.000) 10.4   (±3.495) 3.5   (±3.495) 1.4   (±1.604) 7.6   (±2.660) 2.8   (±3.208) 0.7   (±1.389) 5.3   (±0.684) 3.6171   (±0.170)   2.90
o4-mini 91.0   (±1.389) 75.0   (±0.000) 47.2 (±11.785) 72.9   (±1.389) 50.7   (±7.979) 22.2   (±3.928) 46.5   (±1.389) 19.4   (±6.001) 2.8   (±2.268) 47.5   (±2.063) 1.0908   (±0.115) 12.28
GPT-5-mini 92.4   (±2.660) 83.3   (±7.172) 47.9 (±10.486) 79.2   (±2.778) 59.0   (±3.495) 22.9   (±4.744) 50.7   (±5.727) 22.9   (±1.389) 4.9   (±2.660) 51.5   (±0.154) 0.9598   (±0.096)   9.88
GPT-5 92.4   (±1.389) 91.7   (±0.000) 83.3   (±6.001) 94.4   (±3.928) 83.3   (±3.928) 65.3   (±3.586) 68.8   (±4.167) 47.2   (±3.208) 29.9   (±2.660) 72.9   (±0.812) 0.4525   (±0.075) 19.07
DeepSeek-V3 47.2   (±6.804) 9.7   (±4.811) 2.1   (±2.660) 20.8   (±6.612) 0.7   (±1.389) 0.0   (±0.000) 6.9   (±3.586) 0.0   (±0.000) 0.0   (±0.000) 9.7   (±1.643) 2.8380   (±0.067)   1.59
DeepSeek-R1 89.6   (±3.495) 65.3   (±5.319) 29.2   (±5.319) 66.0   (±9.178) 33.3   (±6.001) 11.1   (±2.268) 38.9   (±8.178) 22.9   (±9.983) 3.5   (±4.167) 40.0   (±2.024) 1.3535   (±0.278) 16.46
QwQ-32b 80.6   (±2.268) 47.2 (±15.713) 15.3   (±4.811) 56.9   (±5.782) 24.3   (±7.979) 0.7   (±1.389) 21.5   (±4.167) 11.1   (±5.072) 0.0   (±0.000) 28.6   (±3.603) 2.4241   (±0.282) 13.99
Qwen-3-235b 82.6   (±4.744) 55.6   (±6.001) 16.7   (±4.536) 54.9   (±9.454) 20.8   (±6.612) 1.4   (±1.604) 31.9   (±7.349) 4.2   (±3.586) 0.7   (±1.389) 29.9   (±1.892) 2.0118   (±0.041) 12.86
Gemini-2.5-flash 90.3   (±3.586) 77.8   (±5.072) 35.4   (±8.893) 84.0   (±5.727) 47.9   (±5.258) 20.1   (±4.744) 47.9   (±4.744) 26.4   (±4.811) 4.9   (±2.660) 48.3   (±1.826) 0.8253   (±0.072) 31.39
Gemini-2.5-pro 98.6   (±1.604) 88.2   (±1.389) 66.7   (±5.072) 90.3   (±4.811) 79.9   (±5.258) 48.6   (±9.488) 61.8   (±2.660) 37.5   (±3.586) 13.9   (±2.268) 65.0   (±2.173) 0.4697   (±0.066) 20.77
@@
GPT-4.1-mini 40.3   (±4.811) 13.9   (±3.928) 4.9   (±2.660) 20.8   (±6.612) 10.4   (±3.495) 4.9   (±2.660) 16.0   (±7.979) 5.6   (±2.268) 0.7   (±1.389) 13.0   (±0.636) 3.0000   (±0.073)   2.96
GPT-4.1 59.7   (±1.604) 18.1   (±3.586) 7.6   (±1.389) 28.5   (±8.295) 9.0   (±2.660) 4.9   (±2.660) 13.2   (±4.744) 2.1   (±2.660) 2.1   (±1.389) 16.1   (±0.729) 2.4598   (±0.233)   4.80
o4-mini 88.9   (±4.536) 69.4   (±6.804) 34.7   (±4.811) 70.1   (±8.599) 45.8   (±3.586) 13.9   (±6.001) 45.8   (±5.319) 15.3   (±2.778) 5.6   (±2.268) 43.3   (±3.422) 1.3706   (±0.099) 11.00
GPT-5-mini 91.7   (±2.268) 72.9   (±4.744) 39.6   (±5.258) 70.1   (±5.727) 44.4   (±7.522) 13.9   (±3.928) 47.9   (±8.893) 18.1   (±4.811) 3.5   (±2.660) 44.7   (±2.069) 1.0295   (±0.058) 10.20
GPT-5 92.4   (±4.167) 93.7   (±2.660) 72.9   (±5.258) 90.3   (±1.604) 77.1   (±1.389) 54.2   (±6.612) 71.5   (±3.495) 43.1   (±3.586) 31.2   (±5.258) 69.6   (±2.055) 0.6133   (±0.152) 18.80
DeepSeek-V3 46.5   (±7.305) 17.4   (±4.167) 4.2   (±1.604) 18.7   (±4.167) 4.2   (±2.778) 0.7   (±1.389) 11.1   (±0.000) 2.1   (±1.389) 2.8   (±2.268) 12.0   (±1.342) 2.7577   (±0.050)   2.50
DeepSeek-R1 77.8   (±6.804) 54.2   (±5.319) 23.6   (±5.782) 60.4   (±3.495) 38.9   (±6.001) 14.6   (±6.159) 38.9   (±5.072) 19.4   (±5.556) 2.1   (±1.389) 36.7   (±2.053) 1.3545   (±0.117) 16.86
QwQ-32b 82.6   (±5.258) 59.7   (±9.213) 23.6   (±3.586) 64.6   (±6.159) 30.6   (±5.072) 2.8   (±0.000) 32.6   (±4.167) 11.1   (±3.928) 0.7   (±1.389) 34.3   (±1.728) 1.7024   (±0.208) 13.84
Qwen-3-235b 87.5   (±4.811) 52.8   (±2.268) 22.9   (±6.564) 64.6   (±4.167) 34.0   (±3.495) 11.1   (±2.268) 29.9   (±4.167) 13.2   (±4.167) 0.0   (±0.000) 35.1   (±1.716) 2.2896   (±0.229) 12.49
Gemini-2.5-flash 85.4   (±5.727) 58.3   (±2.268) 14.6   (±4.167) 70.8   (±9.213) 41.0   (±5.258) 7.6   (±4.167) 40.3   (±4.811) 18.1   (±3.586) 2.1   (±1.389) 37.6   (±1.581) 1.1950   (±0.145) 18.78
Gemini-2.5-pro 95.8   (±1.604) 88.9   (±2.268) 61.1   (±5.072) 88.2   (±2.660) 68.1   (±5.782) 39.6   (±6.944) 63.2   (±6.944) 36.8   (±2.660) 16.0   (±4.167) 62.0   (±0.388) 0.5693   (±0.026) 22.32

### v2

@@ Abstract
Large language models (LLMs) are emerging as powerful tools for scientific law discovery, a foundational challenge in AI-driven science. However, existing benchmarks for this task suffer from a fundamental methodological trilemma, forcing a trade-off between scientific relevance, scalability, and resistance to memorization. Furthermore, they oversimplify discovery as static function fitting, failing to capture the authentic scientific process of uncovering embedded laws through the interactive exploration of complex model systems. To address these critical gaps, we introduce NewtonBench, a benchmark comprising 324 scientific law discovery tasks across 12 physics domains. Our design mitigates the evaluation trilemma by using counterfactual law shifts—systematic alterations of canonical laws—to generate a vast suite of problems that are scalable, scientifically relevant, and memorization-resistant. Moreover, we elevate the evaluation from static function fitting to interactive model discovery, requiring agents to experimentally probe simulated complex systems to uncover hidden principles. Our extensive evaluation of 11 state-of-the-art LLMs reveals a clear but fragile capability for discovery in frontier models: this ability degrades precipitously with increasing system complexity and exhibits extreme sensitivity to observational noise. Notably, we uncover a paradoxical effect of tool assistance: providing a code interpreter can hinder more capable models by inducing a premature shift from exploration to exploitation, causing them to satisfice on suboptimal solutions. These results demonstrate that robust, generalizable discovery in complex, interactive environments remains the core challenge for the future of automated science. By providing a scalable, robust, and scientifically authentic testbed, NewtonBench offers a crucial tool for measuring true progress and guiding the development of next-generation AI agents capable of genuine scientific discovery.
@@ 1 Introduction
Scientific law discovery, the process of distilling complex natural phenomena into concise, predictive mathematical principles, represents a cornerstone of human intellectual achievement, having consistently catalyzed paradigm shifts across the sciences (10.5555/29379; popper2002logic). This endeavor demands a sophisticated interplay of abilities central to the scientific method: formulating hypotheses, methodical experimentation, and synthesizing empirical findings into universal laws (peirce_1958; doi:10.1126/science.1165893).
@@
AI Feynman (udrescu2020aifeynmanphysicsinspiredmethod) 120 Function Partially Tunable Passive Observation
EmpiricalBench (cranmer2023interpretablemachinelearningscience) 9 Function Fixed Passive Observation
@@
LLM-SR (shojaee2025llmsrscientificequationdiscovery) 4 Function Fixed Passive Observation
EvoSLD (lin2025evosldautomatedneuralscaling) 5 Function Fixed Passive Observation
@@
LSR-Transform (shojaee2025llmsrbenchnewbenchmarkscientific) 111 Function Fixed Passive Observation
@@
LSR-Synth (shojaee2025llmsrbenchnewbenchmarkscientific) 128 Function Fixed Passive Observation
PhysSymbol (liu2025mimickingphysicistseyeavlmcentric) 5,000 Function Tunable Passive Observation
@@
The rapid advancement of Large Language Models (LLMs) has brought this grand challenge into sharp focus (openai2024o1preview; comanici2025gemini25pushingfrontier). Exhibiting remarkable emergent capabilities in mathematical reasoning (ahn2024largelanguagemodelsmathematical), logical reasoning (10.1145/3664194), and agentic planning (huang2024understandingplanningllmagents), these models appear to possess foundational skills for scientific inquiry. This has sparked widespread curiosity (Armitage2025; fang2025ainewtonconceptdrivenphysicallaw), crystallized by provocative questions such as whether an LLM could “rediscover Newton’s laws from its own intelligence(he2023future). However, answering such questions requires moving beyond anecdotal observations. To foster and guide progress toward automated science, it is critical to develop a robust testbed to rigorously benchmark the true abilities and inherent limitations of current LLMs in this domain.
@@
While the trilemma highlights the trade-offs between existing paradigms, a more fundamental limitation underlies them all: they frame scientific discovery as a problem of static function discovery, where the goal is simply to fit a mathematical formula to variables in a tabular dataset. This static framing is a stark departure from the authentic scientific process, which relies on the interactive exploration of complex model systems (10.1093/0198247044.001.0001; deSilva2020Discovery)—for instance, J.J. Thomson’s inference of the e/m ratio from cathode ray manipulation. Therefore, to meaningfully benchmark LLMs, evaluations must move beyond static function fitting and embrace the more scientifically authentic challenge of model discovery in interactive environments (fang2025ainewtonconceptdrivenphysicallaw).
To address these limitations and facilitate a more rigorous evaluation, we introduce NewtonBench. Our design is built on two core principles, each targeting a fundamental shortcoming in prior work. First, to resolve the methodological trilemma, we introduce the counterfactual law shift—a technique that systematically alters the mathematical structure of canonical physical laws, for instance, by modifying operators or exponents (see examples in Figure 1). This generates a vast suite of laws that are conceptually grounded but physically novel. This approach simultaneously achieves the scalability of synthetic laws, maintains the scientific relevance of real-world principles, and ensures resistance to memorization by creating problems that cannot be solved by recall, thus forcing models to reason from first principles.
@@
Furthermore, NewtonBench is designed to precisely probe the limits of an LLM’s scientific capabilities. It features two independent dimensions of difficulty control: the intrinsic complexity of the target law (tuned via the counterfactual law shift) and the extrinsic complexity of the surrounding experimental system. This allows for a fine-grained analysis of a model’s breaking points. To isolate the challenge of discovery itself, we introduce an advanced setting where the LLM is provided with a code execution interface for tasks like numerical regression or hypothesis testing. By removing computational limitations as a primary bottleneck, this setup is designed to push models from being merely compute-bound to being truly discovery-bound, revealing the genuine frontiers of their scientific reasoning abilities.
@@ 2 Task Formalization
Figure 1: An illustration of core designs in NewtonBench: experimentation with model system, counterfactual shifts in physical laws from various domains, and agentic exploration settings.
@@ 3.1 Physical Law Curation
Our benchmark’s physical law curation process involves two primary stages: first, the collection of canonical seed laws, and second, the application of Counterfactual Law Shifts to generate novel variations. For the first stage, the selection of our seed laws is guided by three criteria: they must be foundational to their respective fields, span a diverse range of physics domains, and, critically, each must allow for dimensionally consistent mutations (often via adjusting physical constants). The second stage employs Counterfactual Law Shifts—analogous to counterfactual reasoning in physics (Elwenspoek+2012+62+80)—to systematically alter these canonical laws. This approach prevents agents from solving tasks by memorizing known laws, effectively creating a novel hypothetical universe for the agent to discover in each task.
Operationally, these shifts are implemented as mutation operations (Koza1994) on the equation’s expression tree. These mutations can alter either the mathematical operators (e.g., changing an addition to a multiplication) or the values of numeric constants (e.g., modifying an exponent to change a quadratic relationship to a cubic one). A critical consequence of such alterations is the potential disruption of dimensional consistency. To preserve the physical coherence of the modified equation, this disruption is offset by systematically adjusting the dimensional units of the embedded physical constant. This necessity dictates that every target law in our benchmark includes at least one such constant to serve this compensatory role.
Three levels of equation difficulty are categorized based on the cumulative extent of mutations applied to a canonical law [^1] Consequently, equation difficulty arises from two factors: equation complexity and the level of distributional shift.. Easy laws are generated via 1–2 mutations from the original law. Medium laws are subsequently derived from Easy laws through 1–2 additional mutations, and Hard laws are similarly derived from Medium laws. We curated 108 shifted laws from 12 canonical laws under strict guidelines to ensure scientific plausibility, conducted by three domain‑expert co‑authors. The size of our shifted laws is comparable to existing synthetic benchmarks (shojaee2025llmsrbenchnewbenchmarkscientific). Appendix A.3.3 provides the complete set of laws, their mutated forms, and the curation guidelines.
@@ 3.3 Evaluation Metrics
To measure the data fidelity of a submitted law, we adopt the RMSLE metric. Unlike the metric of Normalized Mean Squared Error (NMSE) used in some prior studies (shojaee2025llmsrbenchnewbenchmarkscientific; lin2025evosldautomatedneuralscaling), RMSLE offers superior numerical stability and is better suited for physical quantities that may span multiple orders of magnitude due to its logarithmic scale. Given a dataset of $N$ points with ground-truth values $y_{i}$ and corresponding predictions $\hat{y}_{i}$ from the submitted law, RMSLE is calculated as (implementation details in Appendix A.4.2):
@@
GPT-4.1-mini 20.1   (±1.389) 7.6   (±3.495) 2.8   (±2.268) 8.3   (±4.536) 2.8   (±2.268) 0.0   (±0.000) 4.2   (±1.604) 0.0   (±0.000) 0.0   (±0.000) 5.1   (±0.777) 4.3332   (±0.427) 1.98
GPT-4.1 16.7   (±5.556) 4.2   (±1.604) 1.4   (±1.604) 12.5   (±4.811) 4.9   (±2.660) 2.1   (±2.660) 6.9   (±1.604) 2.8   (±3.208) 0.7   (±1.389) 5.8   (±0.812) 3.8241   (±0.233) 2.82
o4-mini 88.9   (±2.268) 78.5   (±3.495) 52.8   (±9.886) 68.1   (±7.349) 46.5   (±3.495) 22.2   (±2.268) 47.2   (±3.928) 22.9   (±5.258) 2.8   (±2.268) 47.8   (±1.753) 1.2373   (±0.136) 11.57
GPT-5-mini 89.6   (±6.159) 79.9   (±4.167) 60.4   (±4.167) 78.5   (±3.495) 59.0   (±1.389) 26.4   (±4.811) 50.0   (±3.928) 29.2   (±2.778) 4.9   (±2.660) 53.1   (±0.564) 0.7193   (±0.078) 9.41
GPT-5 90.3   (±3.586) 90.3   (±2.778) 87.5   (±6.612) 92.4   (±6.564) 81.9   (±6.612) 64.6   (±2.660) 72.9   (±2.660) 63.2   (±5.727) 40.3   (±5.319) 75.9   (±1.260) 0.2490   (±0.059) 19.15
DeepSeek-V3 39.6   (±1.389) 12.5   (±2.778) 2.1   (±2.660) 18.1   (±4.811) 0.7   (±1.389) 0.0   (±0.000) 6.9   (±2.778) 0.0   (±0.000) 0.0   (±0.000) 8.9   (±1.319) 3.0225   (±0.152) 1.59
DeepSeek-R1 88.2   (±4.167) 72.9   (±3.495) 36.8   (±6.159) 75.0 (±11.785) 41.7   (±5.072) 12.5   (±1.604) 40.3   (±7.349) 20.1 (±10.486) 2.8   (±3.928) 43.4   (±1.369) 1.3839   (±0.172) 16.46
QwQ-32B 74.3   (±5.258) 55.6   (±9.886) 20.8   (±3.586) 47.2   (±7.522) 25.7   (±8.295) 0.7   (±1.389) 20.8   (±5.782) 11.1   (±5.072) 0.0   (±0.000) 28.5   (±3.241) 2.4270   (±0.201) 14.49
Qwen3-235B 81.9   (±1.604) 60.4   (±4.744) 23.6   (±5.782) 51.4   (±5.319) 25.0   (±6.804) 1.4   (±1.604) 28.5   (±8.599) 4.9   (±2.660) 0.7   (±1.389) 30.9   (±0.976) 1.9385   (±0.085) 13.91
Gemini-2.5-flash 93.1   (±2.778) 81.9   (±5.782) 40.3   (±5.782) 79.9   (±6.159) 63.9   (±6.001) 23.6   (±3.586) 41.0   (±5.727) 26.4   (±5.319) 5.6   (±3.208) 50.6   (±1.553) 0.7904   (±0.089) 32.11
Gemini-2.5-pro 96.5   (±2.660) 86.8   (±2.660) 69.4   (±4.536) 86.1   (±3.928) 76.4   (±5.319) 47.2   (±6.804) 62.5   (±3.586) 47.2   (±3.928) 16.7   (±2.268) 65.4   (±1.403) 0.3535   (±0.041) 21.54
@@
GPT-4.1-mini 35.4   (±1.389) 24.3   (±1.389) 16.7   (±2.268) 18.8   (±6.159) 11.1   (±3.928) 6.9   (±4.811) 13.2   (±6.944) 5.6   (±2.268) 0.7   (±1.389) 14.7   (±0.636) 2.7830   (±0.206) 2.92
GPT-4.1 52.1   (±3.495) 31.9   (±1.604) 20.1   (±2.660) 29.2   (±2.778) 10.4   (±4.167) 9.0   (±4.167) 13.2   (±4.167) 2.1   (±2.660) 2.1   (±1.389) 18.9   (±0.772) 2.5044   (±0.229) 4.53
o4-mini 87.5   (±3.586) 70.8   (±4.811) 47.2   (±8.178) 67.4   (±4.744) 52.1   (±4.744) 15.3   (±6.991) 43.1   (±6.991) 17.4   (±2.660) 5.6   (±3.208) 45.1   (±2.038) 1.2078   (±0.077) 10.26
GPT-5-mini 88.2   (±1.389) 73.6   (±4.811) 52.1   (±4.744) 75.0   (±5.072) 51.4   (±7.349) 25.0   (±4.536) 43.8   (±8.599) 20.1   (±3.495) 4.2   (±1.604) 48.1   (±2.153) 0.7412   (±0.111) 9.29
GPT-5 90.3   (±1.604) 89.6   (±2.660) 82.6   (±4.744) 88.2   (±1.389) 78.5   (±4.167) 59.0   (±9.178) 74.3   (±6.159) 59.0   (±4.167) 38.2   (±4.744) 73.3   (±2.384) 0.3793   (±0.117) 18.36
DeepSeek-V3 45.8   (±5.782) 26.4   (±5.782) 9.0   (±2.660) 17.4   (±4.744) 5.6   (±3.928) 1.4   (±1.604) 11.1   (±2.268) 2.1   (±2.660) 0.7   (±1.389) 13.3   (±1.403) 2.8795   (±0.108) 2.23
DeepSeek-R1 73.6   (±4.811) 56.2   (±2.660) 34.7   (±9.755) 58.3   (±3.928) 38.2   (±4.167) 13.2   (±4.167) 38.2   (±1.389) 21.5   (±4.744) 2.8   (±2.268) 37.4   (±1.478) 1.3415   (±0.127) 16.92
QwQ-32B 71.5   (±5.727) 59.0   (±4.744) 25.7   (±2.660) 52.8   (±7.172) 32.6   (±5.727) 2.1   (±1.389) 27.1   (±2.660) 11.1   (±3.928) 0.7   (±1.389) 31.4   (±1.698) 2.0599   (±0.123) 14.62
Qwen3-235B 75.7   (±9.178) 54.2   (±1.604) 27.1   (±4.167) 61.8   (±5.258) 34.7   (±2.778) 10.4   (±2.660) 26.4   (±3.586) 13.2   (±4.167) 0.0   (±0.000) 33.7   (±1.165) 1.9420   (±0.090) 13.30
Gemini-2.5-flash 84.0   (±6.159) 63.9   (±5.072) 28.5   (±7.979) 69.4   (±7.172) 50.0   (±2.268) 13.2   (±4.167) 38.9   (±6.804) 18.1   (±3.586) 2.1   (±1.389) 40.9   (±1.297) 1.1694   (±0.116) 17.73
Gemini-2.5-pro 90.3   (±3.586) 85.4   (±1.389) 66.7   (±2.268) 86.1   (±3.928) 72.9   (±4.167) 43.1   (±5.319) 60.4   (±3.495) 45.8   (±5.782) 22.2   (±3.928) 63.7   (±1.080) 0.3838   (±0.068) 21.93